OpenStudy (anonymous):

Solve for x. Could you also show the steps on how you got the answer. 4^(x+7)=8^(x+3)

OpenStudy (anonymous):

Do you know how to use logarithms?

OpenStudy (anonymous):

Do you know any exponent properties?

OpenStudy (anonymous):

convert both the 4 and the 8 to base 2 so 4 = 2^2 and 8=2^3 now your equation is this 2^2(x+7) = 2^3(x+3) since they have the same base they have the same exponents set the exponents equal to each other 2(x+7) = 3(x+3) factor in the 2 and 3 2x+14=3x+9 combine like terms x= 5

OpenStudy (anonymous):

Ahh, Ok thank you very much I made a foolish error.

OpenStudy (anonymous):

np glad to help